You have a rock with a volume of 15cm3 and a mass of 45 g. What is its density?

  1. You have a rock with a volume of 15cm3 and a mass of 45 g. What is its density?

Grade:10

2 Answers

Debarati
2 years ago
Density= Mass/Volume
Density of the rock= 45g/15cm3= 3 gcm-3 or 3 g/cm3
\thereforeThe density of the rock is 3 g/cm3

I hope you understood.
Nain Singh Rathore
2 years ago
Density= Mass÷Volume.
In the given question,
mass= 45 g
volume= 15cm3
now,
Density= Mass÷Volume= 45÷15=  3g/cm3
Hence, density of the rock is 3g/cm3.